OpenStudy (darkbluechocobo):

Brian needs to protect a science station from a shock wave. Apply the correct equation to determine the energy of the shockwave, if it has a frequency of 8.00 x 10^6 Hz and an amplitude of 3 meters. The mass of the particles is 0.15 kg.

OpenStudy (darkbluechocobo):

@shamim

OpenStudy (shamim):

I m little confused!!!

OpenStudy (darkbluechocobo):

I believe that we are supposed to use

OpenStudy (darkbluechocobo):

@Data_LG2

OpenStudy (anonymous):

hmm... I'm not sure if I understand the problem right you are looking for the energy, which is given by the equation 1/2mv^2 (based from the explanation given) you are also given f= 8 x 10^6 Hz or /s amplitude=3m mass=0.15 kg let me brainstorm a lil bit speed of a wave is given as: v= wavelength x frequency but I don't think that will help because we don't know the wavelength..hmm UNLESS, we use the relationship between E and A, which is given as: E=1/2kA^2 (derived when Ek=0) hmm.. give me a sec to figure this one out @_@

OpenStudy (anonymous):

can you post a pic of the whole question including the equation that you posted?

OpenStudy (darkbluechocobo):

oh i found that in my reading s:

OpenStudy (darkbluechocobo):

that is the whole question, i didnt get it cause I have nothing bout shockwaves T_T

OpenStudy (anonymous):

Here's what I came up with-really 100% NOT sure http://sketchtoy.com/64665986 draw tool is not working, so I used that ^

OpenStudy (darkbluechocobo):

does it say 2amg?

OpenStudy (anonymous):

yeah, but don't follow it..i'm sure it's wrong.. i'm still finding a more logical way